OpenStudy (anonymous):

Ill give medal for good answer? Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (amistre64):

well, giving a medal for a good answer is simply a means of thanking people, its not really a motivation towards blurbing out an answer tho

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

if you let us know where it is in this that your having issues at, we could help.

OpenStudy (amistre64):

otherwise the question is designed to get youto think for yourself

OpenStudy (anonymous):

prettymuch the whole thing

OpenStudy (amistre64):

Create a unique example of dividing a polynomial by a monomial and provide the simplified form. do you know the what a poly and a mono are? that would be a start Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

sort of it has to do with the number of exponents and variables right?

OpenStudy (amistre64):

yes; moreso, its has to do with the number of terms in an equation (how many add and subtracts there are)

OpenStudy (anonymous):

can u give me an example?

OpenStudy (amistre64):

so, poly means many; and mono means 1 we want a many termed equation divided by a single termed equation to start out with

OpenStudy (amistre64):

3x + 4y -7z = 9 is a poly; theres alot of addition and subtraction

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

5abc is a mono in that there is absolutely no addition or subtraction going on

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

so for this question; give me your favorite poly and your favorite mono :)

OpenStudy (anonymous):

i guess a mono could be 7ab^2 + 5c - 4b^2?

OpenStudy (anonymous):

i think its a mono

OpenStudy (amistre64):

thats a poly; since there is alot of adding and subtracting going on between terms. but we can use it

OpenStudy (anonymous):

ok and amono could be 6ab?

OpenStudy (amistre64):

perfect :)

OpenStudy (amistre64):

to create the division of poly/mono we just: \[\frac{ 7ab^2 + 5c - 4b^2}{6ab}\]

OpenStudy (anonymous):

then it said i need the simplified form and then two ways used to simplifty

OpenStudy (amistre64):

yeah, that part is a bit vague to me; the simplified form might be when we split the top and reduce tho ....

OpenStudy (anonymous):

then a way that i would check my quotient for accuracy

OpenStudy (anonymous):

dang dude how much u typing...

OpenStudy (amistre64):

\[\frac{ 7ab^2 + 5c - 4b^2}{6ab}\to \ \frac{ 7ab^2}{6ab} + \frac{5c}{6ab} - \frac{4b^2}{6ab}\] \[\frac{ 7abb}{6ab} + \frac{5c}{6ab} - \frac{4bb}{6ab}\] \[\frac{ 7\cancel{a}b\cancel{b}}{6\cancel{a}\cancel{b}} + \frac{5c}{6ab} - \frac{4b\cancel{b}}{6a\cancel{b}}\] \[\frac{7b}{6} + \frac{5c}{6ab} - \frac{4b}{6a}\]

OpenStudy (amistre64):

the latex coding is tedious, but makes it prettier :)

OpenStudy (anonymous):

and the bottom one is the simplified version?

OpenStudy (amistre64):

yes, with any luck I typed it up in order

OpenStudy (amistre64):

except for 4/6 = 2/3

OpenStudy (anonymous):

wait so whats the answer in typing not the latex?

OpenStudy (amistre64):

but thats reducing, which might mean simplifying too

OpenStudy (amistre64):

in typing its a mess ....

OpenStudy (amistre64):

7b/6 + 5c/6ab -2a/3b maybe

OpenStudy (anonymous):

the thing im doing the final product doesnt have any / in it though

OpenStudy (anonymous):

it has answers like 2x^2y^3 + 3y

OpenStudy (amistre64):

the issue here is that you have material to go by while I have to take a blind stab at it and hope for the best :)

OpenStudy (amistre64):

that result there would depend on what they used for poly and mono

OpenStudy (amistre64):

your results will differ

OpenStudy (amistre64):

one possibility might be in that we do long division on it instead of splitting the numerator

OpenStudy (anonymous):

can u help me on last part cause it says i need a way to check it?

OpenStudy (anonymous):

latex again?

OpenStudy (amistre64):

7b/6 + 5c/6ab -2b/3a -------------------- 6ab ) 7ab^2 +5c -4b^2 (7ab^2) ------------- 0+ 5c (5c) ----------- 0 - 4b^2 (-4b^2) --------- 0 so yeah, the results are the same

OpenStudy (amistre64):

no, just trying to get the ascii in a sensible format lol

OpenStudy (anonymous):

i think thats it for that question...

OpenStudy (amistre64):

yep, to check you just multiply it back thru to see that it comes back to the original again

OpenStudy (anonymous):

alright well thanks for the help :)

OpenStudy (amistre64):

yw, and good luck